Q2) Sales are a function of advertising in newspapers and magazines (X, Y).              10

S = 100X0.5 Y0.5

Price of advertising in newspapers and magazines are Rs.30 and Rs.40 respectively. The total budget for advertising is Rs.1200. For maximizing the sales, find out the best combination of advertisement in newspapers and magazines by using lagrangian multiplier


1
Expert's answer
2021-06-07T11:30:57-0400
"f(x,y)=100x^{0.5}y^{0.5}-\lambda(30x+40y-1200)"

"\frac{\delta f}{\delta x}=50(\frac{y}{x})^{0.5}-30\lambda"


"\frac{\delta f}{\delta y}=50(\frac{x}{y})^{0.5}-40\lambda"


"y=0.75x"

"x=20"

"y=15"